OpenStudy (anonymous):

The area of a parallelogram is A = bh. Solve for h (height) in this formula.

OpenStudy (anonymous):

h=A/b

OpenStudy (anonymous):

h=Ab h=A/b h=b/a h=A-b

OpenStudy (anonymous):

A = bh Isolate h by dividing both sides by b. h = a/b

OpenStudy (anonymous):

Second one

OpenStudy (anonymous):

Ok thanks @jcpd910 @datusernametho

OpenStudy (anonymous):

No problem. Datusernametho shouldn't have given you the answer without an explanation, but he/she is new.